OpenStudy (anonymous):

is this problem solvable via integration by parts? (sin(x))/(x^2+1)

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ \sin x }{ x^2 + 1} = ?\]

OpenStudy (anonymous):

i presume that u = sin (x) dv = \[\frac{ 1 }{ x^2+1 }dx\]

OpenStudy (anonymous):

or is this not solvable at all by that method? it tends to give a zero eh...

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ \sin x }{ x^2+1 } ={ \sin x \times \tanh^{-1} x } + \int\limits_{}^{} \tanh^{-1}x \times \cos x\]

OpenStudy (lgbasallote):

tanh?

OpenStudy (lgbasallote):

\[\int \frac{dx}{x^2 + 1} = \tan^{-1} x\] right?

OpenStudy (anonymous):

yes sir.

OpenStudy (anonymous):

oh sorry. it should be tan -1

OpenStudy (lgbasallote):

right

OpenStudy (anonymous):

let me rewrite \[\int\limits_{}^{}\frac{ \sin x }{ x^2 + 1} = \sin x \times \tan^{-1} x - \int\limits_{}^{} \tan^{-1} x \times \cos x\]

OpenStudy (lgbasallote):

right.. now integrate by parts again

OpenStudy (anonymous):

then lhs=rhs...i mean...same integrand will come on RHS

OpenStudy (anonymous):

dv = \[\frac{ 1 }{ x^2 + 1 }\] u = cos x so, \[\int\limits_{}^{}\frac{ \sin x }{ x^2 + 1 } = \sin x \times \tan^{-1}x - (\cos x \times \tan^{-1} + \int\limits_{}^{} \tan^{-1} x \times \sin x)\] is that it akash, or i had it wrong?

OpenStudy (lgbasallote):

should be \[\int \frac{\sin x}{x^2 + 1}\] because it's vdu

OpenStudy (anonymous):

sorry if i'm a bit confused... hahaha

OpenStudy (lgbasallote):

dv = tan^-1 x v = 1/x^2 + 1 right?

OpenStudy (lgbasallote):

or no?

OpenStudy (anonymous):

oh yeah. you're right LGS. i was confused. hahahahaha.

OpenStudy (anonymous):

i remembered i isolated it because of LIATE. =)))

OpenStudy (lgbasallote):

y =tan^-1 x tan y = x sec^2 y = 1 y = 1/sec^2 y y = 1/x^2 + 1 yup that's right

OpenStudy (lgbasallote):

so you should have \[\int \frac{\sin x}{x^2 + 1} = \sin x \tan^{-1} x - \frac{\cos x}{x^2 + 1} - \int \frac{\sin x}{x^2 + 1}\] right?

OpenStudy (anonymous):

oh... and then move the integral on the other side, divide by to and eat your bacon? =)))

OpenStudy (lgbasallote):

remember to eat the fats first

OpenStudy (anonymous):

yessir. but, wolfram alpha has a weird answer. =))) ((-1 + E^2)*CosIntegral[I - x] + (-1 + E^2)*CosIntegral[I + x] + I*(1 + E^2)*(SinIntegral[I - x] + SinIntegral[I + x]))/(4*E)

OpenStudy (lgbasallote):

huh

OpenStudy (lgbasallote):

that's interesting...

OpenStudy (anonymous):

it has imaginary numbers. that looked like some f-d up s-it to me. =))

OpenStudy (lgbasallote):

\[\int \frac{\sin x}{x^2 + 1}\] \[u = \frac{1}{x^2 + 1} \implies du = -\frac{2x}{(x^2 + 1)^2}\] \[dv = \sin x \implies v = -\cos x\] \[\int \frac{\sin x}{x^2 + 1} = -\frac{\cos x}{x^2 + 1} - 2\int \frac{x\cos x}{x^2 + 1}\] yes.. i think we made a mistake a while ago...this is complex...very complex

OpenStudy (lgbasallote):

i saw our mistake earlier...we made wrong substitution in the second IBP

OpenStudy (anonymous):

ah yes. we forgot to apply LIATE immediately. instead of LIATE, ETAIL happened. =)))

OpenStudy (lgbasallote):

well..that wasnt the problem i was talking about...

OpenStudy (lgbasallote):

order of LIATE doesn't really matter

OpenStudy (lgbasallote):

however.... you let \[u = \sin x \implies du = \cos x\] \[dv = \frac{1}{x^2 + 1} \implies v = \tan^{-1}x\] so... \[\int \frac{\sin x}{x^2 + 1} = \sin x \tan^{-1} - \int \tan^{-1} \cos x dx\] see th mistake?

OpenStudy (lgbasallote):

what we did earlier was + instead of -

OpenStudy (anonymous):

yeap. that should have been a negative. =)))

OpenStudy (lgbasallote):

yes. so now, this is equally difficult and complex

OpenStudy (anonymous):

unfortunately there is No closed form for this integral in terms of elementary functions

OpenStudy (lgbasallote):

indeed

OpenStudy (lgbasallote):

sorry to have dragged it. i thought there was hope

OpenStudy (anonymous):

no problem. My professor and I have been working on it for a few days already. Looks like it needs something called, "taylor", is it? =)))

OpenStudy (lgbasallote):

i hear it's a very swift method